OpenStudy (anonymous):

May you please help me? What is the slope of a line that passes through the point (−1, 1) and is parallel to a line that passes through (3, 6) and (1, −2)?

OpenStudy (anonymous):

the first point isnt needed......it can be solved by finding the slope of line joining 2nd and 3rd points........ the formula for slope is M=(y2-y1)/x2-x1 M=(-2-6)/(1-3) M=-8/-2 M=4